Unfortunately, arsenic occurs naturally in some groundwater. A mean arsenic level of \mu=8.0 parts per billion (ppb) is considered safe for agricultural use. A well in Texas us used to water cotton crops. This well is tested regular basis for arsenic. A random sample of 37 tests gave a sample mean \bar{x}=7.2 pbb of arsenic, with s = 1.9 pbb. Does this information indicate that the mean of arsenic in this well is less than 8 pbb? Use \alpha=0.01


1
Expert's answer
2021-12-14T08:40:11-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq 8"

"H_1:\mu<8"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=36" degrees of freedom, and the critical value for a left-tailed test is "t_c = -2.434494."

The rejection region for this left-tailed test is "R = \{t: t < -2.434494\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{7.2-8}{1.9/\sqrt{37}}\approx -2.561163"

Since it is observed that "t = -2.561163 <-2.434494= t_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=36" degrees of freedom, "t=-2.561163" is "p=0.007384," and since "p=0.007384<0.01=\alpha,"  it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than "8," at the "\alpha = 0.01"significance level.